# Breaking Focus: Contextual Distraction Curse in Large Language Models

### v1

@@ 5 Conclusion
In this work, we identify a critical weakness, termed Contextual Distraction Vulnerability (CDV), where models struggle with irrelevant but semantically coherent context. Using an efficient tree-based search method, we generate CDV examples across four datasets, causing a significant performance drop in the state-of-the-art models. Mitigation strategies, particularly post-targeted training, show promise in improving robustness. This highlights CDV as an ability-level challenge, emphasizing the need to address this issue to ensure model reliability.
Appendix A Experiment Details
@@ A.1 Experiment Settings
Experimental details of baseline methods. To validate the effectiveness of our tree-based search framework, we implemented two baseline perturbation approaches for comparison. The Elaborated method performed semantic-preserving length augmentation by expanding original questions with explanatory clauses and redundant contextual information while maintaining core semantics. The Prompt-only baseline utilized our perturbation prompt template (details in Appendix D) through Claude-3.5-Sonnet for automatic disturbance generation without subsequent search optimization. For a fair comparison, all baseline methods processed the same 100 original questions from four datasets using Claude-3.5-Sonnet as the executor. The enhanced questions were evaluated under identical zero-shot CoT settings across seven target models. This demonstrates the crucial role of our tree-based search mechanism in identifying optimal perturbation combinations rather than relying on simple length expansion or single-pass prompt perturbations. The results of this experiment are summarized in Figure 3.
@@
Experimental details of classifier. We used the 1200 original questions from the scale-up experiment, splitting them into training, test, and validation sets. Specifically, 80 percent of the data was allocated to training, with 10 percent of the training set reserved for validation, and the remaining 10 percent was used for testing. For the prompt-based classifiers, we designed specific prompts to guide the models in determining whether a problem was hard to perturb. We evaluated the classification performance of seven models: GPT-4o-mini, GPT-4o, Llama-3.1-8B, Gemma-2-27B, Gemma-2-2B, Qwen2.5-1.5B, and Qwen2.5-7B. A baseline configuration without any classifier was also included for comparison. The effectiveness of these classifiers was measured using the F1-score with beta equal to 0.5, which prioritizes precision over recall. For the training-based classifiers, we used supervised fine-tuning with LoRA on four open-source models: Llama-3.1-8B, Gemma-2-2B, Qwen2.5-1.5B, and Qwen2.5-7B. The training was conducted on a single RTX 4090 GPU, with a learning rate set to 1e-4 and a total of five epochs. The performance of these fine-tuned classifiers was also evaluated using the F1-score on the test set. This experimental design allowed us to compare the utility of prompt-based and training-based classifiers in identifying hard-to-perturb questions. The results of this experiment are summarized in Figure 4, Table 2, Figure 5 and Table 7.
@@
Table 7: Performance of the classifier under Prompt-Based and Fine-Tuned methods. The table reports Precision, Recall, and F0.5 scores for both Prompt-Based (left) and Fine-Tuned (right) classifiers. Fine-tuned models are marked in the Fine-Tuned columns. The baseline represents the performance of the system without using the classifier.
@@
Experimental details of mitigating CDV. We curated approximately 1200 preference data pairs. Each preference pair consisted of a question, a correct answer, and an incorrect answer collected from model responses in prior experiments. To ensure a fair evaluation, we guaranteed that enhanced questions originating from the same original question did not appear in both the training and test sets. The data was split into training, validation, and test sets, with 80 percent of the data used for training, 10 percent of the training set reserved for validation, and 20 percent allocated to testing. For prompt-based enhancement, we designed new prompt templates aimed at improving model focus on the core question content and tested them on seven models: GPT-4o-mini, Gemma-2-27B, Llama-3.1-8B, Qwen2.5-72B, o1-mini, GPT-4o, and Claude-3.5-Sonnet. For training-based enhancement, we fine-tuned three open-source models, namely Gemma-2-2B, Qwen2.5-7B, and Phi-3.5-mini. Using the Direct Preference Optimization algorithm, the fine-tuning was performed on two RTX 4090 GPUs with a learning rate set to 2e-4 and five epochs. The preference loss was implemented with a sigmoid activation function. The fine-tuned models were evaluated against three high-performance baseline models, specifically GPT-4o, GPT-4o-mini, and Qwen2.5-72B, using the original zero-shot with CoT prompting templates on the test set. This experiment provided insights into the effectiveness of both prompt-based and training-based approaches in improving model robustness against enhanced questions. The results of this experiment are summarized in Table 4 and Table 5.
A.2 Experiment Analysis
@@
Appendix B Human Evaluation
@@
Appendix C Case Study
Figures 11, 12, 13, 14, 15, 16, and 17 showcase the specific response performances of various models when confronted with both original and enhanced questions.
From Figures 8, 9 and 10, we present cases illustrating the changes in responses to enhanced questions by the Gemma-2-2B, Phi-3.5-mini, and Qwen2.5-7B following training-based improvements.
@@ Appendix C Case Study
Appendix D Prompt Template
[…]
@@ Appendix D Prompt Template
[…]
@@
[…]

### v2

@@ 4 Conclusion
In this work, we propose a framework to assess the contextual robustness of language models by generating adaptive distractions, which are semantically coherent but task-irrelevant additions. Our tree-based search method produces challenging examples that induce consistent performance drops across models and datasets. Among mitigation strategies, post-training methods such as DPO offer the most reliable improvements. Ultimately, our approach offers a scalable tool for evaluating and improving LLM reliability in real-world applications. Future work will integrate our distraction generation into training loops to further strengthen contextual robustness.
Appendix A Related Work
A.1 Contextual Robustness in LLMs
Recent studies have highlighted that LLMs often fail when presented with semantically coherent yet task-irrelevant contextual information. Shi et al. [5] introduced a benchmark demonstrating that irrelevant context can severely degrade LLM accuracy in arithmetic reasoning tasks. Liu et al. [6] showed that model performance significantly drops when key information is placed in the middle of long contexts, indicating positional sensitivity in attention mechanisms. Similar findings have emerged in narrative distraction scenarios [39], misprimed probe studies [40], and irrelevant document retrieval contexts [41]. Collectively, these works underscore contextual distraction as a prevalent vulnerability affecting modern LLMs.
A.2 Generation of Contextual Perturbations
To systematically probe LLM sensitivity to irrelevant context, various methods have been developed to generate targeted perturbations. Zhu et al. [33] proposed a dynamic evaluation approach using meta probing agents that restructure tasks to surface latent weaknesses in model behavior. Similarly, optimization-based prompt injection methods have been explored to create adversarial inputs aimed at exploiting model biases or alignment issues [42, 43]. Chatziveroglou et al. [39] further validate that even semantically coherent but irrelevant narratives can significantly reduce LLM accuracy. Despite these efforts, existing perturbation generation techniques typically focus on altering prompts or instructions without necessarily preserving answer correctness or targeting semantic coherence explicitly.
A.3 Mitigation Strategies
Several approaches have attempted to address the issue of contextual distraction by enhancing model robustness [44]. PromptBreeder [45] employs evolutionary strategies to optimize task prompts, implicitly strengthening model robustness against perturbations. Wu et al. [46] introduced prompting strategies instructing models to explicitly ignore irrelevant information, and Wang et al. [38] showed that self-consistency decoding can improve reliability by aggregating predictions from multiple reasoning paths. Zhu et al. [47] explored attention mechanisms, identifying internal attention directions to guide models toward more relevant context. Moreover, Yoran et al. [41] demonstrated the effectiveness of fine-tuning on mixed-relevance data to improve robustness in retrieval-augmented scenarios. Direct preference optimization (DPO) [48], a targeted fine-tuning technique, has gained particular attention for its effectiveness in enhancing model alignment and resilience to distracting inputs. While prompt-based approaches remain attractive due to their simplicity, our findings indicate they have limited efficacy against adaptive distractions. Our experiments highlight the stronger robustness achieved by fine-tuning strategies, validating the effectiveness of targeted mitigation approaches in addressing contextual distraction.
Appendix B Experiment Details
@@ B.1 Experiment Settings
Experimental details of baseline methods. To validate the effectiveness of our tree-based search framework, we implemented two baseline perturbation approaches for comparison. The Irrelevant Context Augmentation (ICA) method performed semantic-preserving length augmentation by expanding original questions with task-irrelevant but semantically coherent contextual information, such as explanatory clauses or redundant details. The Single-Prompt Distraction (SPD) baseline utilized our perturbation prompt template (details in Appendix G) through Claude-3.5-Sonnet for automatic distraction generation without subsequent search optimization. For a fair comparison, all baseline methods processed the same 100 original questions from four datasets using Claude-3.5-Sonnet as the executor. The enhanced questions were evaluated under identical zero-shot CoT settings across seven target models. This demonstrates the crucial role of our tree-based search mechanism in identifying optimal perturbation combinations rather than relying on simple length expansion or single-pass prompt perturbations. The results of this experiment are summarized in Table 2.
@@
Experimental details of classifier. We used 1200 original questions from 4 datasets, splitting them into training, test, and validation sets. Specifically, 80 percent of the data was allocated to training, with 10 percent of the training set reserved for validation, and the remaining 10 percent was used for testing. For the prompt-based classifiers, we designed specific prompts to guide the models in determining whether a problem was hard to perturb. We evaluated the classification performance of seven models: GPT-4o-mini, GPT-4o, Llama-3.1-8B, Gemma-2-27B, Gemma-2-2B, Qwen2.5-1.5B, and Qwen2.5-7B. A baseline configuration without any classifier was also included for comparison. The effectiveness of these classifiers was measured using the F1-score with beta equal to 0.5, which prioritizes precision over recall. For the training-based classifiers, we used supervised fine-tuning with LoRA on four open-source models: Llama-3.1-8B, Gemma-2-2B, Qwen2.5-1.5B, and Qwen2.5-7B. The training was conducted on a single RTX 4090 GPU, with a learning rate set to 1e-4 and a total of five epochs. The performance of these fine-tuned classifiers was also evaluated using the F1-score on the test set. This experimental design allowed us to compare the utility of prompt-based and training-based classifiers in identifying hard-to-perturb questions. The results of this experiment are summarized in Figure 4, Table 3, Figure 3 and Table 8.
To evaluate the tradeoff between precision and recall in our classifier analysis, we report the Fβ score with $\beta=0.5$. This metric places greater emphasis on precision, which is desirable in our use case. The score is defined as:
$\text{F}_{\beta}=(1+\beta^{2})\times\frac{\text{Precision}\times\text{Recall}}{(\beta^{2}\times\text{Precision})+\text{Recall}}.$ (15)
@@
Table 8: Performance of the classifier under Prompt-Based and Fine-Tuned methods. The table reports Precision, Recall, and F0.5 scores for both Prompt-Based (left) and Fine-Tuned (right) classifiers. Fine-Tuned models are marked in the Fine-Tuned columns. Baseline represents the performance of the system without using classifier.
@@
Experimental details of mitigation. We curated approximately 1200 preference data pairs. Each preference pair consisted of a question, a correct answer, and an incorrect answer collected from model responses in prior experiments. To ensure a fair evaluation, we guaranteed that enhanced questions originating from the same original question did not appear in both the training and test sets. The data was split into training, validation, and test sets, with 80 percent of the data used for training, 10 percent of the training set reserved for validation, and 20 percent allocated to testing. For prompt-based enhancement, we designed new prompt templates aimed at improving model focus on the core question content and tested them on seven models: GPT-4o-mini, Gemma-2-27B, Llama-3.1-8B, Qwen2.5-72B, o1-mini, GPT-4o, and Claude-3.5-Sonnet. For training-based enhancement, we fine-tuned three open-source models, namely Gemma-2-2B, Qwen2.5-7B, and Phi-3.5-mini. Using the Direct Preference Optimization algorithm, the fine-tuning was performed on two RTX 4090 GPUs with a learning rate set to 2e-4 and five epochs. The preference loss was implemented with a sigmoid activation function. The fine-tuned models were evaluated against three high-performance baseline models, specifically GPT-4o, GPT-4o-mini, and Qwen2.5-72B, using the original zero-shot with CoT prompting templates on the test set. This experiment provided insights into the effectiveness of both prompt-based and training-based approaches in improving model robustness against enhanced questions. The results of this experiment are summarized in Table 5 and Table 6.
B.2 Preliminary Experimental Results
To support our claim in the introduction regarding the limited effectiveness of static distraction methods, we present a comparison of performance drops induced by our adaptive distraction framework and a representative static method, GSM-IC [5], in Table 9. The results demonstrate that while GSM-IC causes minimal performance drops (average 1.8%) on advanced models, our method achieves significantly larger drops (average 45%), highlighting its potency in challenging LLM contextual robustness.
Table 9: Comparison of performance drops (%) on GSM-IC [5] and our adaptive distraction.
Model GPT-4o-mini GPT-4o Qwen2.5-72B Gemma-2-27B Claude-3.5-sonnet o1-mini
GSM-IC [5] 4.1 1.2 1.4 2.2 0.8 1.2
B.3 Experiment Analysis
@@
B.4 Cross-Dataset Generalization of the Classifier
We present additional evaluation of the classifier’s ability to generalize across datasets. As shown in Table 10, the $F_{0.5}$ score remains stable even when the classifier is tested on a domain it was not trained on.
Table 10: $F_{0.5}$ scores of the classifier trained on MMLU, CommonsenseQA, and OpenbookQA, and tested on the held-out TruthfulQA dataset.
Model Orig. Test Cross-Dataset
Gemma-2-2B 0.724 0.735 (+0.011)
Qwen2.5-1.5B 0.712 0.698 (-0.014)
Llama-3.1-8B 0.816 0.782 (-0.034)
B.5 Additional Prompting Strategies
We further evaluate widely used prompting methods on the same set of adaptively distracted questions. Despite using more sophisticated prompting, the performance gains are marginal, confirming that contextual distraction remains a persistent challenge even under varied prompting schemes.
Table 11: Accuracy of various prompting strategies on adaptive distraction examples. All numbers are averaged across four datasets. Values in parentheses indicate drop from clean accuracy.
Model Vanilla Zero +CoT [37] Few +CoT [37] Few-shot (ICL) [37] Self-Consistency [38]
GPT-4o-mini 0.185 (-0.705) 0.211 (-0.679) 0.238 (-0.652) 0.262 (-0.628) 0.272 (-0.618)
GPT-4o 0.390 (-0.550) 0.391 (-0.549) 0.455 (-0.485) 0.443 (-0.497) 0.461 (-0.479)
Qwen2.5-72B 0.334 (-0.486) 0.343 (-0.477) 0.349 (-0.471) 0.344 (-0.476) 0.352 (-0.468)
Llama-3.1-8B 0.247 (-0.420) 0.251 (-0.416) 0.272 (-0.395) 0.277 (-0.390) 0.285 (-0.382)
B.6 Evaluation on MATH-500 Reasoning Benchmark
To test whether adaptive distraction applies beyond factual QA, we extend our evaluation to MATH-500, a benchmark composed of free-form math questions with clear ground truth. To integrate this dataset into our framework, we prompt a strong LLM (Claude-3.5-Sonnet) to generate three plausible but incorrect answer options for each original question, converting the problem into a multiple-choice format. As shown in Table 12, the majority of models suffer substantial performance drops, demonstrating that our method remains effective even in formal reasoning domains.
Table 12: Performance on MATH-500 before and after adaptive distraction. Values in parentheses denote accuracy drops.
Model Original w/ Adaptive Distraction
GPT-4o-mini 0.805 0.302 (-0.503)
GPT-4o 0.813 0.488 (-0.325)
Gemma-2-27B 0.658 0.242 (-0.416)
Llama-3.1-8B 0.650 0.252 (-0.398)
Qwen2.5-72B 0.871 0.527 (-0.344)
Claude-3.5-Sonnet 0.828 0.516 (-0.312)
o1-mini 0.958 0.857 (-0.101)
Interestingly, o1-mini shows notable resilience, but even strong reasoning models are not immune. These results confirm that adaptive distraction reveals a broader attention failure affecting all tasks with structured ground truth, not just factual QA.
B.7 Comparing Supervised Fine-Tuning and DPO
We compare DPO-based mitigation with standard SFT to evaluate whether the performance gains from DPO are significantly higher than those from SFT, and whether these gains come at the cost of performance on the original, unperturbed examples.
Effectiveness on Adaptive Distraction.
Table 13 shows that while SFT offers mild improvements over the base model, DPO achieves substantially higher accuracy under adaptive distraction (AD).
Table 13: Comparison of SFT and DPO on adaptive distraction (AD). Original (AD) refers to the base model’s accuracy on perturbed inputs without mitigation.
Model Original (AD) SFT (AD) DPO (AD)
Gemma-2-2B 0.257 0.305 (+0.048) 0.432 (+0.175)
Qwen2.5-7B 0.212 0.278 (+0.066) 0.440 (+0.228)
Phi-3.5-mini 0.195 0.261 (+0.066) 0.680 (+0.485)
Impact on Clean Accuracy.
We further assess whether these robustness gains come at the cost of performance on clean inputs. Table 14 shows that DPO-tuned models maintain nearly all of their original accuracy, indicating minimal performance trade-off.
Table 14: Accuracy on clean and adaptive distraction (AD) inputs before and after DPO fine-tuning. Clean (Orig) refers to the base model’s accuracy on original inputs without distraction.
Model Clean (Orig) AD (Orig) Clean (DPO) AD (DPO)
Gemma-2-2B 0.450 0.257 0.398 0.432 (+0.034)
Qwen2.5-7B 0.480 0.212 0.411 0.440 (+0.029)
Phi-3.5-mini 0.720 0.195 0.697 0.680 (-0.017)
These results confirm that DPO provides substantial robustness gains under adaptive distraction, while preserving performance on clean questions. In contrast, SFT yields only modest improvements and does not fully address the distraction vulnerability.
Appendix C Human Evaluation
@@
Appendix D Overall Algorithm
We show the overall algorithm in Algorithm 1.
Appendix E Limitations and Broader Impacts
While our adaptive distraction generation framework provides valuable insights into LLM contextual robustness, its long-term efficacy must be considered within the rapidly evolving landscape of LLM development, necessitating continuous adaptation of such probing methodologies. Furthermore, our current focus on semantically coherent, task-irrelevant contextual additions, while demonstrably effective, represents one facet of potential distractions; future work could explore a broader taxonomy of disturbances and extend generalization to a wider array of complex tasks and domains. Crucially, as with any potent diagnostic tool, the responsible development and deployment of such adaptive probing techniques are paramount to ensure they contribute positively to LLM safety and trustworthiness, mitigating risks of misuse and fostering a more robust LLM ecosystem.
Output: Candidate problem list $L$
Initialize priority queue $\mathcal{Q}\leftarrow\emptyset$ and candidate list $L\leftarrow\emptyset$;
foreach $P=\langle Q,A_{\text{gt}},\mathcal{A}_{\text{inc}}\rangle\in D$ do
 if $p(y=1\mid Q)=C(Q)<\tau_{C}$ then
    continue // Filter low-potential questions using classifier ;
   end if
 if $r_{M}(P)$ = 0 then
      Add $P$ to $L$;
    continue
   end if
   Add root node $P$ to $\mathcal{Q}$;
end foreach
while $\mathcal{Q}\neq\emptyset$ do
   Pop $P^{\prime}=\arg\max_{P\in\mathcal{Q}}\mathcal{V}(P)$, $\mathcal{Q}\leftarrow\mathcal{Q}\setminus\{P^{\prime}\}$;
  Generate $k=|\mathcal{A}_{\text{inc}}|$ child nodes for $P^{\prime}$ using $P_{\text{proxy}}$;
 for each child node $P^{\prime}_{j}$ do
      Compute semantic shift $S(P,P^{\prime}_{j})$ and length ratio $\text{len}(P^{\prime}_{j})/\text{len}(P)$ // Semantic shift check and computing length ratio ;
    if $S(P,P^{\prime}_{j})=1$ and $\text{len}(P^{\prime}_{j})/\text{len}(P)\leq\lambda$ then
         Compute value $\mathcal{V}(P^{\prime}_{j})$;
      end if
    else
         Discard $P^{\prime}_{j}$;
      end if
   end for
 if $|\mathcal{C}_{0}(P^{\prime})|>n_{1}$ then
      Add top $n_{1}$ nodes from $\mathcal{C}_{0}(P^{\prime})$ to $L$
     terminate branch;
    continue;
   end if
 if $r_{M}(P^{\prime}_{j})=1,\,\forall P^{\prime}_{j}$ then
    continue // Skip nodes where all children are unpromising ;
   end if
  Add $P^{\prime}_{j}$ nodes to $\mathcal{Q}$;
 if $m_{i+1}>m_{i},\,\forall i\in\{1,2,\dots,l-1\}$ then
    continue // Prune monotonically increasing success rate branches ;
   end if
end while
return $L$;
Algorithm 1 Overall Algorithm
Appendix F Case Study
Figures 10, 11, 12, 13, 14, 15, and 16 showcase the specific response performances of various models when confronted with both original and enhanced questions.
From Figures 7, 8 and 9, we present cases illustrating the changes in responses to enhanced questions by the Gemma-2-2B, Phi-3.5-mini, and Qwen2.5-7B following training-based improvements.
@@ Appendix F Case Study
Appendix G Prompt Template
[Instruction]
@@ Appendix G Prompt Template
Please carefully read the question below and provide a solution from the choices. You must choose the model’s final answer from one of the choices. Let’s think step by step!
@@
Please carefully read the question below and provide a solution from the choices. You must choose the model’s final answer from one of the choices. Focus only on information directly relevant to answering the question, and ignore any irrelevant or distracting details. Let’s think step by step!